a goldsmith combines an alloy that equals 4.20 per ounce with an alloy that costs 1.70 per ounce. How many ounces of each were used to make a mixture of 200 ozcosting 2.45
Answer by ankor@dixie-net.com(22740)

a goldsmith combines an alloy that equals 4.20 per ounce with an alloy that costs 1.70 per ounce.
How many ounces of each were used to make a mixture of 200 ozcosting 2.45
:
let x = amt of $4.20 alloy
then since the result is to be 200 oz,
200-x = amt of $1.70 oz stuff
:
4.2x + 1.7(200-x) = 2.45(200)
4.2x + 340 - 1.7x = 490
4.2x - 1.7x = 490 - 340
2.5x = 150
x = 150/2.5
x = 60 oz of the $4.20 alloy
and
200 - 60 = 140 oz of the $1.70 alloy